Question #3

Georgia Clay and Gravel was updating the MPS record for one of its products, Smell Fresh Cat Litter.

  1. Complete the following MPS time-phased record.

Smell Fresh

On Hand

Week

1

2

3

4

5

6

7

8

Forecast

20

20

20

30

30

30

30

30

Orders

5

3

2

Projected available balance

20

50

30

10

30

50

20

40

10

Available to promise

MPS

50

50

50

50

MPS Lot Size = 50

  1. The following events occurred in week 1:
  • Actual demand during week 1 was 10 units.
  • Marketing forecasted that 50 units would be needed for week 9.
  • An order for 12 in week 2 was accepted.
  • An order for 8 in week 3 was accepted.
  • An order for 25 in week 4 was accepted.
  • The MPS in week 1 was produced as planned.

Update the record below after rolling through time.

Smell Fresh

On Hand

Week

2

3

4

5

6

7

8

9

Forecast

Orders

Projected available balance

60

(Hint)

Available to promise

MPS

MPS Lot Size = 50
Answer #1

(a)

Smell Fresh On Hand Week
1 2 3 4 5 6 7 8
Forecast 20 20 20 30 30 30 30 30
Orders 5 3 2 0 0 0 0 0
Projected Avl Bal. 20 50 30 10 30 50 20 40 10
Available to promise 60 50 50 50
MPS 50 50 50 50

Note:

ATP for the first period = Beginning Inventory + MPS of the first period - Sum of orders up to the next MPS.

ATP for other periods = MPS of that period - Sum of orders up to the next MPS.

(b)

Smell Fresh On Hand
2 3 4 5 6 7 8
Forecast 20 20 30 30 30 30 30
Orders 15 10 25 0 0 0 0
Projected Avl Bal. 60 40 20 40 60 30 50 20
Available to promise 35 25 50 50
MPS 50 50 50

B 2 Smell Fresh 20 30 30 Forecast Orders Projected Avl Bal. Available to promise MPS On Hand 2 5 8 20 30 30 30 =3+12 =2+8 =20